Find the LCM and HCF of the following integers by applying the prime factorisation method.  17, 23 and 29 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Writing the product of prime factors for all the three numbers, we get,

17=17×1
23=23×1
29=29×1
Therefore,

HCF(17,23,29) = 1
LCM(17,23,29) = 17 × 23 × 29 = 11339

Thanks